Japan’s combination of X and Y

Canada’s combination of X and Y

100x and 0y

50x and 0y

50x and 100y

25x and 75y

0x and 200y

0x and 150y

  1. What are the opportunity costs for 1 unit of X and Y for Japan?
  2. What are the opportunity costs for 1 unit of X and Y for Canada?
  3. What is the comparative advantage in producing for Japan?
  4. What is the comparative advantage in producing for Canada?
Answer #1

A)

InJapan

To produce 50 X units from 0 to 50, give up 100 units of Y

So for 1 X = give up 2 units of Y

So opportunity cost of​​​​​ production of X = 2 units of Y

opportunity cost of production of Y = .5 units of X

B) Canada

to produce 25 units of X, give up (150-75) = 75 Y

For 1 X, give up 75/25 = 3 Y

So opportunity cost of X = 3 units of Y

opportunity cost of Y = 1/3 units of X

c) nation with lower opportunity cost of production of a good (in terms of other good) has comparative advantage in that good

thus Japan has comparative advantage in production of X ,

Bcoz opportunity cost of production of X is lower in Japan

( One needs to give up only 2 y for one X in Japan , as against 3 Y in Canada )

D) canada has comparative advantage in production of Y